OpenStudy (anonymous):

WILL MEDAL please solve 8x = 4x - 32 8 4 -4 -8

OpenStudy (haseeb96):

8x-4x=-32 4x =-32 x=-32/4 x= -8 this is your correct answer

OpenStudy (paki):

just compare the similar terms and solve... @camilacabelllo